Solve the inequality. x/3 - x-1/2 > 1 there is a line under the greater than sign. Thank you!

Answer by checkley79(3341)

x/3 - x-1/2 >= 1
-2x/3>=1+1/2
-2x/3>=1.5 cross multiply
-2x>=1.5*3
-2x>=4.5 divide by -2 (note the sign change)
x<=4.5/-2
x<=-2.25 ans.